OpenStudy (anonymous):

How do I solve -4x+5y=0 3x+2y=23 by substitution?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

Is it necessary by substitution? U can multiply equation (-4x+5y=0) by 2 and the other equation by 5 and then substract the two. Or you can use equation(-4x+5y=0) and put y= 4x/5 and put it in the second equation

OpenStudy (anonymous):

Method 1: 1. Multiply Eqn 1 by 2 -8x+10y = 0 2. Multiply Eqn 2 by 5 15x + 10y = 115 3. Subtract these 2 equations - 23x = -115 x = 5 4. Substitute x = 5 in Eqn 1

OpenStudy (anonymous):

Thank You so much for your help. :)